OpenStudy (anonymous):

A + B <=> C + 2D. One mole of A and one mole of B are placed in a .400L container. After equilibrium has been established what is the Kc?

OpenStudy (anonymous):

*After equilibrium has been established there is .2 moles of C.

OpenStudy (anonymous):

Using the molar ratios of the balanced equation and knowing that there are .2 moles of C AFTER equilibrium is established we know the following: Moles of A = .2 Moles of B = .2 Moles of D = .4 I'm going to assume this is a gas. Therefore we can find the concentrations of all the species using \[mol/L =M\] Therefore the concentrations are as follows: A= .5 B= .5 c= .5 D = 1 Now that we know the concentrations we can make an equilibrium expression (products/ reactants) \[Kc = \left[ A \right]\left[ B \right]/\left[ C \right]\left[ D \right]^{2}\] Plug in the concentrations where they belong and solve for Kc you should get Kc=.5